OpenStudy (anonymous):

I REALLY NEED SOMEONE TO HELP ME WITH AT LEAST ONE OF THESE PROBLEMS.. PLEASE The solid is similar to a larger solid with the given scale factor, find the surface area and volume of the larger solid. 1) Cylinder: scale factor 1:2, SA=208pi in^2, V=320pi in^3

OpenStudy (anonymous):

The first thing that needs to be done is find the radius and height of the original cylinder. Using \[SA = 2\pi r h = 208 \pi\] and \[V = \pi r^2 h = 320 \pi\] we get: \[(1)r h = 104\] \[(2)r^2 h = 320\] Solving for h in equation (1), we get \[h = 104/r\] so substituting this into equation (2), we get \[r = 320/104\] and by using the formula we got for h, we find \[h = 104^2/320\] Since the scale is 1:2, I'm asuming that this means the radius and height are doubled. Hence our new r and h are \[r = 320/52 = 80/13\] \[h=104^2/160=10816/160=338/5\] Now by the surface area formula, \[SA = 2* \pi *80/13*338/5 = 832 \pi \] By the Volume formula, \[V = \pi * r^2 * h = \pi * (80/13)^2 * 338/5 = 2560 \pi \] I think that's it.

OpenStudy (anonymous):

thank you.. can you look at my other two please i have been doing these for a week and cant figure them out